If the sampled population has a mean of 48 and standard deviation of 16, then the mean and the standard deviation for the sampling distribution of x-bar for n = 16 are

  

4 and 1

   

12 and 4

48 and 4

   

48 and 1

   

48 and 16

I know the answer already. please explain!
Answer #1

Solution :

Given that,

The sampling distribution of mean is ,

\mu\bar x = 48

The sampling distribution of standard deviation is ,  

\sigma\bar x = \sigma / \sqrt n = 16 / \sqrt 16 = 16 / 4 = 4

48 and 4